### v3

# Orbital Optical Raman Lattice

Abstract
Spin and orbital are two basic degrees of freedom that play significant roles in exploring exotic quantum phases in optical lattices with synthetic spin-orbit coupling (SOC) and high orbital bands, respectively. Here, we combine these two crucial ingredients for the first time by proposing a completely new orbital optical Raman lattice scheme to explore exotic high-orbital Bose condensates with Raman-induced SOC in a square lattice. We find that both the SOC and $p$-orbital interactions influence the condensed state of bosons. Their interplay results in two novel high-orbital many-body quantum phases: the uniform angular momentum superfluid phase, which exhibits a global topological chiral orbital current characterized by a uniform Chern number, and the two-dimensional topological spin-orbital supersolid phase, which is characterized by the spin and orbital angular momentum density wave patterns and topological excitations with opposite Chern numbers, respectively protecting the chiral and antichiral edge modes in the neighboring supersolid clusters. Our scheme may open a new avenue for exploring exotic SOC and high-orbital physics in optical lattices, and is expected to advance the experimental realization of novel supersolids in higher dimensions.
preprint: APS/123-QED
Introduction.
Ultracold atomic gases provide ideal platforms for quantum simulation due to their pristine nature and full controllability Bloch et al. (2008); Gross and Bloch (2017). Among them, investigations into synthetic gauge fields Jaksch and Zoller (2003); Juzeliūnas and Öhberg (2004); Osterloh et al. (2005); Ruseckas et al. (2005); Liu et al. (2006); Dalibard et al. (2011); Goldman et al. (2014); Chen et al. (2018); Aidelsburger et al. (2022) and spin-orbit couplings (SOCs) Liu et al. (2009); Juzeliūnas et al. (2010); Lin et al. (2011); Campbell et al. (2011); Sau et al. (2011); Wang et al. (2012); Cheuk et al. (2012); Anderson et al. (2012); Galitski and Spielman (2013); Liu et al. (2014); Zhai (2015); Wu et al. (2016); Burdick et al. (2016); Song et al. (2016); Chen et al. (2016); Kolkowitz et al. (2017); Zhang and Liu (2018); Wang et al. (2018) have attracted widespread interests. Over the past decade, experimental realizations of novel SOCs in one-dimension Lin et al. (2011); Wang et al. (2012); Cheuk et al. (2012); Burdick et al. (2016); Song et al. (2016); Kolkowitz et al. (2017), two-dimension Wu et al. (2016); Huang et al. (2016); Meng et al. (2016); Sun et al. (2018a, b), and three-dimension Wang et al. (2021a) have enabled the simulation of various exotic topological models with cold atoms, such as topological semimetals Wang et al. (2021a), quantum anomalous Hall insulator Liang et al. (2023), and non-Hermitian topological phases Zhao et al. (2023). Additionally, extending synthetic SOC to strongly correlated regime facilitates the realization of non-Abelian dynamical gauge fields Zhou et al. (2023). Besides simulating topological quantum phases Goldman et al. (2010); Béri and Cooper (2011); Liu et al. (2013); Jotzu et al. (2014); Wang et al. (2014); Zhang and Zhou (2017); Song et al. (2018); Yi et al. (2019), the quest for supersolids characterized by diagonal and off-diagonal long-range order in ultracold atoms has never ceased Boninsegni and Prokof’ev (2012); Ritsch et al. (2013); Recati and Stringari (2023). Currently, the supersolid phase has been observed in long-range interacting dipolar quantum gases Li et al. (2017); Böttcher et al. (2019); Tanzi et al. (2019); Chomaz et al. (2019); Norcia et al. (2021); Bland et al. (2022); Casotti et al. (2024) and optical cavities Léonard et al. (2017a, b).
High-orbital (e.g. $p$ and $d$) systems in optical lattices also received considerable attention due to their rich degrees of freedom, which can give rise to exotic orbital physics Zhao and Liu (2008); Liu et al. (2010); Cai et al. (2011); Lewenstein and Liu (2011); Li et al. (2013); Sowiński et al. (2013); Dutta et al. (2015); Li and Liu (2016); Kock et al. (2016). With the orbital degree of freedom, one can not only simulate the behavior of electrons in realistic materials but also, more crucially, uncover new concepts and phenomena that have no prior analog in electronic systems Li and Liu (2016), such as the high-orbital Bose-Einstein condensates (BECs) with novel orbital ordering Isacsson and Girvin (2005); Kuklov (2006); Liu and Wu (2006); Wu et al. (2006); Larson et al. (2009); Wirth et al. (2011); Li et al. (2012); Xu et al. (2016); Di Liberto et al. (2016); Wang and Liu (2017); Li et al. (2018); Niu et al. (2018); Pan et al. (2020); Jin et al. (2021); Wang et al. (2021b); Huang et al. (2022); Wang et al. (2023). To date, significant progress have been made in the study of $p$-orbital BEC in optical lattices. For instance, the $p$-orbital BEC in a square lattice manifests as a superfluid phase with a staggered angular momentum order that breaks time-reversal symmetry Liu and Wu (2006). In a triangular lattice, it results in staggered loop current orders Wu et al. (2006); Wang et al. (2023), and in a $sp^{2}$-orbital hexagonal lattice, it exhibits Potts-nematic superfluid order Jin et al. (2021) or atomic chiral superfluidity with topological excitations Wang et al. (2021b). Such phases reveal the remarkable richness of higher orbital coherence whether or not breaking the time-reversal symmetry.
Figure 1: (a) Generic setup for orbital optical Raman lattice. The bias magnetic field $\bm{B}$ is parallel to the x direction. A pair of standing waves, $\bm{E}_{xy}$ and $\bm{E}_{yz}$, produce a square lattice. Moreover, the Raman field can be achieved by applying another plane wave $\bm{E}_{zx}$. (b) Optical-dipole transition diagram for Raman coupling in cold bosons (${}^{87}\mathrm{Rb}$) coupled to two pairs of laser beams ($\bm{E}_{xy}$, $\bm{E}_{zx}$) and ($\bm{E}_{yz}$, $\bm{E}_{zx}$). Here the hyperfine states |1,1|1,-1$\rangle$| 1 , - 1 ⟩ and |1,0|1,0$\rangle$| 1 , 0 ⟩ of ${}^{87}\mathrm{Rb}$ are used to emulate the ground states $|g_{\uparrow,\downarrow}\rangle$. (c) The sketch for the nearest-neighbor spin-conserved hopping for $s$ and $p_{x,y}$ orbitals and the spin-flipped hopping between $s$ and $p_{x,y}$. Here, only the $p_{x}$ orbital is displayed, since the hopping terms for $p_{x}$ and $p_{y}$ are related by $C_{4}$ symmetry. (d) Single-particle spectrum. The $C_{4}$-related points $\bm{k}_{0}^{1}$, $\bm{k}_{0}^{2}$, $\bm{k}_{0}^{3}$, $\bm{k}_{0}^{4}$ are the lowest energy points that lie on the diagonal lines in the Brillouin zone. Thereinafter, $V_{0}=5.0E_{r}$, $M_{0}=1.0E_{r}$ (tight-binding parameters: $t_{s}=0.0658E_{r}$, $t_{p}^{\parallel}=0.4228E_{r}$, $t_{p}^{\perp}=0.0658E_{r}$, $t_{so}=0.1059E_{r}$, $\mu_{s}=5.1909E_{r}$, and $\mu_{p}=5.9049E_{r}$), $m_{z}=0.0173E_{r}$. Besides, $\bm{k}_{0}^{1}=(0.154,0.154)\pi$. $E_{r}$ is the recoil energy.
In this letter, we propose an orbital optical Raman lattice scheme that integrates the synthetic SOC Wu et al. (2016); Liu et al. (2014); Wang et al. (2018) and $p$-orbital condensates Isacsson and Girvin (2005); Kuklov (2006); Liu and Wu (2006), predicting two exotic quantum many-body phases with nontrivial interaction-induced topology in the square lattice: the uniform angular momentum superfluid (UAMSF) and the 2D topological spin-orbital supersolid (SOSS), both of which go beyond the scope of previous studies on synthetic SOC and high-orbital physics. We show that these condensed states of bosons are governed by the nontrivial interplay effects between on-site $p$-orbital interactions and Raman-induced SOC, surpassing the limitations of pure orbital lattice gas models without synthetic SOC. The UAMSF exhibits a uniform angular momentum order, supporting uniform chiral topological excitations. In contrast, the 2D topological SOSS shows staggered spin and angular momentum cluster patterns, leading to the chiral and antichiral topological edge excitations in the neighboring supersolid clusters. Our work offers a new perspective for the combined study of nontrivial synthetic SOC and high-orbital physics, which were previously considered separate fields in optical lattice research, and will promote the realization of novel supersolids without long-range interactions in ultracold atom systems.
Model.
We start with the Hamiltonian for ultracold bosons trapped in an orbital optical Raman lattice, given by $H_{0}=\frac{p_{x}^{2}}{2m}+\frac{p_{y}^{2}}{2m}+V(x,y)+M(x,y)\sigma_{x}+m_{z}\sigma_{z}$. Here, $V(x,y)=-V_{0}[\cos^{2}(\frac{\pi}{a}x)+\cos^{2}(\frac{\pi}{a}y)]$ and $M(x,y)=-M_{0}[\cos(\frac{\pi}{a}x)+\cos(\frac{\pi}{a}y)]$ represent the normal lattice potential and Raman field, respectively. The Pauli matrices $\bm{\sigma}$ act on the subspace spanned by $|s\rangle\otimes|g_{\uparrow}\rangle$ and $|p\rangle\otimes|g_{\downarrow}\rangle$, where the spin $|g_{\uparrow,\downarrow}\rangle$ is defined from hyperfine ground states, $|s\rangle$ and $|p\rangle$ represent local orbital states. The model Hamiltonian $H_{0}$ can be realized with high feasibility, through the three-laser configuration shown in Fig. 1(a). Two lasers form standing waves, $\bm{E}_{xy}=\hat{y}E_{0}\cos(k_{0}x)$ and $\bm{E}_{yz}=\hat{z}E_{0}\cos(k_{0}y)$, linearly polarized along $y$ and $z$ axes, respectively, and propagating in $x$-$y$ plane. The third laser forms a plane wave, $\bm{E}_{zx}=\hat{x}E_{z}e^{-\mathrm{i}k_{z}z}$ propagating along $z$ direction and is $x$-polarized. The standing waves $\bm{E}_{xy}$ and $\bm{E}_{yz}$ coupling to excited states with a red detuning $\Delta$ contribute to the diagonal square optical lattice potentials $V(x,y)=\hbar(|\bm{E}_{xy}|^{2}/\Delta+|\bm{E}_{yz}|^{2}/\Delta)$. Additionally, $\bm{E}_{xy}$, $\bm{E}_{yz}$, and $\bm{E}_{zx}$ induce the Raman field components $M(x)$ and $M(y)$ via two two-photon transitions as illustrated in a $\bm{\varLambda}$-type configuration [Fig.1(b)]. In experiment one can easily set that the Raman field only couples $|s\rangle\otimes|g_{\uparrow}\rangle$ and $|p\rangle\otimes|g_{\downarrow}\rangle$ by putting such two states be nearly resonant for the two-photon transitions except for a tunable two-photon detuning $\delta$, which defines an effective Zeeman splitting $m_{z}=\hbar\delta/2$, while all other orbital states are far detuned SM .
With the above implementation scheme, bosons can occupy the spin-orbital locking states $\phi_{s_{\uparrow}}$, $\phi_{p_{x\downarrow}}$, and $\phi_{p_{y\downarrow}}$. As shown in Fig.1(c), the lattice potential $V(x,y)$ governs spin-conserved nearest-neighbour hopping ($t_{s}$ and $t_{p}$), while the Raman field $M(x,y)$ contributes to spin-flipped nearest-neighbour hopping ($t_{so}$). Since the Raman field $M(x,y)$ has twice the period of the optical lattice potential $V(x,y)$ and is symmetric with respect to each lattice site center of $V(x,y)$, the spin-flipped hopping is staggered in the $\nu$ ($\nu=x,y$) direction with $t_{so}^{(\bm{i},\bm{i}\pm\bm{e}_{v})}=\pm(-1)^{i_{v}}t_{\mathrm{SO}}$ SM . The staggered factor $(-1)^{i_{\nu}}$ represents a nontrivial momentum transfer between spin-up and spin-down states in the Raman coupling, and can be absorbed by the transformation $b_{\bm{i},p_{\nu},\downarrow}\rightarrow-(-1)^{i_{\nu}}b_{\bm{i},p_{\nu},\downarrow}$, yielding the tight-binding Hamiltonian
$\hat{H}_{0}=$ $(\mu_{s}+m_{z})\sum_{\bm{i}}n_{\bm{i},s,\uparrow}+(\mu_{p}-m_{z})\sum_{\bm{i},\nu}n_{\bm{i},p_{\nu},\downarrow}$
-- $t_{s}\sum_{\langle\bm{i},\bm{j}\rangle}b_{\bm{i},s,\uparrow}^{\dagger}b_{\bm{j},s,\uparrow}-\sum_{\bm{i},\mu,\nu}(\bar{t}_{p}^{\mu\nu}b_{\bm{i},p_{\mu},\downarrow}^{\dagger}b_{\bm{i}+\bm{e}_{\nu},p_{\mu},\downarrow}+\mathrm{H.c.})$
$+$ $t_{\mathrm{SO}}\sum_{\bm{i},\nu}\sum_{\delta_{\nu}=\pm\bm{e}_{v}}\Big{[}\text{sgn}(\delta_{\nu})b_{\bm{i},p_{\nu},\downarrow}^{\dagger}b_{\bm{i}+\delta_{\nu},s,\uparrow}+\mathrm{H}.\mathrm{c.}\Big{]},$ (1)
where $b_{\bm{i},l,\sigma}^{\dagger}$ ($b_{\bm{i},l,\sigma}$) ($l=s,p_{x},p_{y}$) denotes the creation (annihilation) operator of boson. $\bar{t}_{p}^{\mu\nu}=t_{p}^{\parallel}\delta_{\mu\nu}+t_{p}^{\perp}(1-\delta_{\mu\nu})$ ($\mu$, $\nu$=$x$, $y$), $n_{\bm{i},l,\sigma}=b_{\bm{i},l,\sigma}^{\dagger}b_{\bm{i},l,\sigma}$, and $\mu_{s,p}$ denote the on-site energies. A key feature is that the momentum transfer by Raman field forces the band minima of $\epsilon_{p_{x}\downarrow}(\bm{k})$ ($\epsilon_{p_{y}\downarrow}(\bm{k})$) to move from $\bm{Q}_{x}=(\pi,0)$ ($\bm{Q}_{y}=(0,\pi)$) to $\bm{\Gamma}$ SM . Further, the remaining 2D SOC effect can result in a single-particle spectrum with four $C_{4}$-connected minima at $\bm{k}_{0}^{1}$, $\bm{k}_{0}^{2}$, $\bm{k}_{0}^{3}$, and $\bm{k}_{0}^{4}$ [Fig.1(d)], leading to the major novel physics predicted in the present orbital optical Raman lattice. A quadratic band touch point protected by time-reversal symmetry exists at $\bm{\Gamma}$ Sun et al. (2012).
Figure 2: (a) Schematic of the variational calculation. Here, $C_{4}$-connected points $\bm{k}_{0}^{n}$, $\bm{k}_{c}^{n}$, and $\tilde{\bm{k}}_{c}^{n}$ ($n=1\sim 4$) represent the energy minima of the single-particle spectrum, the variational momenta within the areas surrounded by red circles, and the condensation momenta along the blue diagonal lines, respectively. (b) Ground state phase diagram. $m_{z}$ and $n_{0}$ denote the effective Zeeman splitting and boson condensation density, respectively. There are three types of ground states: $C_{4}$-symmetry-broken superfluid phase ($C_{4}$-SBSF), two-dimensional spin-orbital supersolid phase (2D SOSS), and uniform angular momentum superfluid phase (UAMSF). Parameters: $V_{0}=5.0E_{r}$, $M_{0}=1.0E_{r}$, and $g=0.01E_{r}$ ($U_{s}=0.1044E_{r}$, $U_{sp}=0.0428E_{r}$, $U_{p}=0.0618E_{r}$, and $\widetilde{U}_{p}=0.0175E_{r}$).
The total Hamiltonian includes also the contact interaction $V(\bm{r}-\bm{r}^{\prime})=g_{\sigma\sigma^{\prime}}\delta(\bm{r}-\bm{r}^{\prime})$ ($\sigma,\sigma^{\prime}=\uparrow,\downarrow$), whose form in the lattice model reads SM
$\hat{H}_{\mathrm{int}}=$ $\sum_{\bm{i}}\Big{[}\frac{U_{s}}{2}n_{\bm{i},s,\uparrow}(n_{\bm{i},s,\uparrow}-1)+U_{sp}n_{\bm{i},s,\uparrow}\sum_{\nu}n_{\bm{i},p_{\nu},\downarrow}$
$+$ $\frac{U_{p}}{2}\sum_{\nu}n_{\bm{i},p_{\nu},\downarrow}(n_{\bm{i},p_{\nu},\downarrow}-1)+2\widetilde{U}_{p}n_{\bm{i},p_{x},\downarrow}n_{\bm{i},p_{y},\downarrow}$
$+$ $\frac{\widetilde{U}_{p}}{2}(b_{\bm{i},p_{x},\downarrow}^{\dagger}b_{\bm{i},p_{x},\downarrow}^{\dagger}b_{\bm{i},p_{y},\downarrow}b_{\bm{i},p_{y},\downarrow}+\mathrm{H.c.})\Big{]},$ (2)
where $U_{s}=g_{\uparrow\uparrow}\int\mathrm{d}\bm{r}|\phi_{s,\uparrow}(\bm{r}-\bm{r}_{\bm{i}})|^{4}$, and the coefficients $U_{sp}$, $U_{p}$, and $\widetilde{U}_{p}$ can also be similarly defined with $g_{\sigma\sigma^{\prime}}\approx g$. Note that $U_{p}$ refers to the intra-orbital interaction for the $p_{\nu\downarrow}$ ($\nu=x,y$) orbitals, while $\widetilde{U}_{p}$ refers to the inter-orbital interaction between the $p_{x\downarrow}$ and $p_{y\downarrow}$ orbitals.
Ground state ansatz and phase diagram.
Usually, the non-interacting BEC will take place at the single-particle band minimum points $\bm{k}_{0}^{n}$ ($n=1\sim 4$). However, since these lowest energy points of Hamiltonian $\hat{H}_{\mathrm{0}}$ do not generally correspond to the energy minima of the total Hamiltonian $\hat{H}_{\mathrm{0}}+\hat{H}_{\mathrm{int}}$, the interaction changes the populations in the $s_{\uparrow}$, $p_{x\downarrow}$, and $p_{y\downarrow}$ orbitals, which shifts the condensation momenta around $\bm{k}_{0}^{n}$ ($n=1\sim 4$). Under weak interaction, the condensed ground state ansatz can be given by $|g\rangle\sim e^{\sqrt{N_{0}}b^{\dagger}}|\mathrm{vac}\rangle$, where $b^{\dagger}$, expressed as $b^{\dagger}=\sum_{n=1}^{4}\sum_{l_{\sigma}}\gamma_{\bm{k}_{c}^{n}}\beta_{\bm{k}_{c}^{n},l_{\sigma}}b_{\bm{k}_{c}^{n},l_{\sigma}}^{\dagger}$ ($l_{\sigma}=s_{\uparrow},p_{x\downarrow},p_{y\downarrow}$), represents the quasi-particle creation operator and $N_{0}$ is the particle number of BEC. Here, $C_{4}$-related points $\bm{k}_{c}^{n}$ [Fig.2(a)] denote the variational momenta, with the variational parameters $\gamma_{\bm{k}_{c}^{n}}$ and $\beta_{\bm{k}_{c}^{n}}$ satisfying $\sum_{n}|\gamma_{\bm{k}_{c}^{n}}|^{2}=1$ and $\sum_{l_{\sigma}}|\beta_{\bm{k}_{c}^{n},l_{\sigma}}|^{2}=1$, respectively. Note that $|\gamma_{\bm{k}_{c}^{n}}|^{2}$ represents the condensation distribution probability of bosons at $\bm{k}_{c}^{n}$, and $|\beta_{\bm{k}_{c}^{n},l_{\sigma}}|^{2}$ denotes the orbital population probability. To identify the condensation momenta $\tilde{\bm{k}}_{c}^{n}$ and the condensation parameters ($\gamma_{\tilde{\bm{k}}_{c}^{n}}$, $\beta_{\tilde{\bm{k}}_{c}^{n},l_{\sigma}}$), we search for the minimum of the energy density functional $\langle g|\hat{\mathcal{H}}_{0}+\hat{\mathcal{H}}_{\mathrm{int}}|g\rangle/N$ by simulated annealing algorithm, with $N$ the number of lattice sites [see Supplementary for details SM ]. As depicted in Fig.2(a), $\tilde{\bm{k}}_{c}^{n}$ ($n=1\sim 4$) lie on the blue diagonal lines and move toward $\bm{\Gamma}$ (or coincide with it), and are connected by $C_{4}$ symmetry.
Next, our calculation finds that the interplay between the $p$-orbital interaction in Eq.(2) and the Raman-induced 2D SOC is highly nontrivial and leads to the emergence of exotic orbital orders with distinct novel topology in the condensate state $|G\rangle$. The $p$-orbital interaction in the present orbital optical Raman lattice tends to generate uniform non-zero orbital angular momentum (OAM) order $\langle G|L_{i,z}|G\rangle$ ($L_{i,z}=-\mathrm{i}b_{\bm{i},p_{x},\downarrow}^{\dagger}b_{\bm{i},p_{y},\downarrow}+\mathrm{H.c.}$) Liu and Wu (2006); SM to reduce the interaction energy, leading to $\beta_{\tilde{\bm{k}}_{c}^{n},p_{x\downarrow}}=\mathrm{\pm i}\beta_{\tilde{\bm{k}}_{c}^{n},p_{y\downarrow}}$. However, the Raman-induced 2D SOC in Eq.(1) prefers $\beta_{\tilde{\bm{k}}_{c}^{n},p_{x\downarrow}}=\pm\beta_{\tilde{\bm{k}}_{c}^{n},p_{y\downarrow}}$, suppressing the uniform OAM $\langle L_{i,z}\rangle$. Thus, it is this competitive relationship between them that gives rise to three distinct phases: (1) When the SOC dominates over the $p$-orbital interaction, the condensate will occur at one of the $\tilde{\bm{k}}_{c}^{n}$ ($n=1\sim 4$) with $\beta_{\tilde{\bm{k}}_{c},p_{x\downarrow}}=\pm\beta_{\tilde{\bm{k}}_{c},p_{y\downarrow}}$, resulting in a $C_{4}$-symmetry-broken superfluid phase without OAM order; (2) In contrast, if the $p$-orbital interaction plays the leading role, the bosons condense at the $\bm{\Gamma}$ point with $\beta_{\bm{\Gamma},p_{x\downarrow}}=\mathrm{-i}\beta_{\bm{\Gamma},p_{y\downarrow}}$, and the condensate exhibits uniform OAM order; (3) Importantly, in the intermediate regime, where the $p$-orbital interaction and SOC are comparably strong, the bosons condense equally at the $\tilde{\bm{k}}_{c}^{n}$ ($n=1$sim 4), with the orbital parameters given by $\beta_{\tilde{\bm{k}}_{c}^{n},p_{x\downarrow}}=-(-1)^{n}\beta_{\tilde{\bm{k}}_{c}^{n},p_{y\downarrow}}$, rendering the total system a novel OAM density-wave pattern known as the SOSS phase. Fig.2(b) shows the phase diagram plotted versus condensate density $n_{0}$ and $m_{z}$, which govern the $p$-orbital interaction. A most important feature of the phases in (2) and (3) [orange and green areas in Fig.2(b)] is that they host different types of topological quasiparticles, whose emergence necessitates both the $p$-orbital interaction and the 2D SOC. We examine them below in detail.
Uniform angular momentum superfluid.
Figure 3: (a) OAM order $\langle L_{i,z}\rangle$ and (b) topological quasiparticle excitation spectrum for UAMSF. The parameters: $m_{z}=0.08E_{r}$, $n_{0}=2.0$. Other parameters are taken as the same as those in Fig.2.
Adjust the Zeeman splitting $m_{z}$ such that onsite energy $\epsilon_{p\downarrow}(\bm{\mathrm{\Gamma}})$ is well below $\epsilon_{s\uparrow}(\bm{\mathrm{\Gamma}})$. With the dominant $p$-orbital population, the $p$-orbital interaction governs the ground state. Consequently, bosons condense at the $\bm{\mathrm{\Gamma}}$ with orbital parameters $\beta_{\bm{\mathrm{\Gamma}},s_{\uparrow}}=0$ and $\beta_{\bm{\mathrm{\Gamma}},p_{x\downarrow}}=\mathrm{-i}\beta_{\bm{\mathrm{\Gamma}},p_{y\downarrow}}$. As shown in Fig.3(a), this ground state exhibits a uniform OAM order $\langle L_{\bm{i},z}\rangle=2n_{0}|\beta_{\bm{\mathrm{\Gamma}},p_{y\downarrow}}|^{2}$ in real space, rendering the UAMSF which breaks time-reversal symmetry. While the ground state is dominated by $p$-orbital interaction, the quasiparticle excitations are actually governed by both the $p$-orbital interaction and the 2D SOC, and exhibit nontrivial topology characterized by Chern numbers.
Figure 4: (a) Spatial distribution of $\langle L_{\bm{i},z}\rangle$ for 2D SOSS phase. The positive ($\langle L_{\bm{i},z}\rangle>0$) and negative ($\langle L_{\bm{i},z}\rangle<0$) OAM clusters are staggered, with $\sum_{\mathrm{Global}}\langle L_{\bm{i},z}\rangle=0$. Inset: the relative phase difference $\theta_{i}=\text{Im}\Big{[}\mathrm{ln}\langle b_{\bm{i},p_{x},\downarrow}^{\dagger}b_{\bm{i},p_{y},\downarrow}\rangle\Big{]}$ between the $p_{x}$ and $p_{y}$ orbitals along path $\bf{r}$. (b) Topological edge excitations of the positive A and negative B clusters. The arrows represent the flow directions of the edge currents. (c) Spin field $(\langle\sigma_{\bm{i},x}\rangle,\langle\sigma_{\bm{i},y}\rangle)$ displays a topological pattern (skyrmion). (d) Spin component $\langle\sigma_{\bm{i},z}\rangle$ presents another staggered high-and low-density clusters pattern. Here, parameters are set as $m_{z}=0.06E_{r}$, $n_{0}=2.0$. Besides, the condensation momenta are $\widetilde{\bm{k}}_{c}^{1}=(k_{c},k_{c})$, $\widetilde{\bm{k}}_{c}^{2}=(-k_{c},k_{c})$, $\widetilde{\bm{k}}_{c}^{3}=(-k_{c},-k_{c})$, and $\widetilde{\bm{k}}_{c}^{4}=(k_{c},-k_{c})$, where $k_{c}=\pi/10$.
In order to obtain the quasiparticle excitation spectrum, here we employ the Bogoliubov theory to derive the Bogoliubov-de Gennes (BdG) Hamiltonian ^$\hat{\mathcal{H}}_{\mathrm{BdG}}=\frac{1}{2}\sum_{\bm{k}}\Psi_{\bm{k}}^{\dagger}\mathcal{H}_{\mathrm{BdG}}\Psi_{\bm{k}}$ SM , where $\Psi_{\bm{k}}^{\dagger}=(b_{\bm{\Gamma}+\bm{k},l,\sigma}^{\dagger},b_{\bm{\Gamma}-\bm{k},l,\sigma})$ denotes the Nambu basis. Performing the Bogoliubov transformation, we have $T_{\bm{k}}^{\dagger}\mathcal{H}_{\mathrm{BdG}}(\bm{k})T_{\bm{k}}=E_{\bm{k}}$. Here para-unitary matrix $T_{\bm{k}}$ satisfies $T_{\bm{k}}^{\dagger}\tau_{z}T_{\bm{k}}=\tau_{z}$ ($\tau_{z}=\sigma_{z}\otimes\mathrm{\bm{I}}_{3\times 3}$) and the diagonal terms of $E_{\bm{k}}$ represent the excitation spectrum Shindou et al. (2013). As shown in Fig.3 (b), an interaction-induced topological gap opens at the $\bm{\Gamma}$ point Sun et al. (2012) and, together with the 2D SOC, separates the first band from the second and third bands, leading to the Chern number of the first band $\mathrm{Ch}^{1}=1$, while the total topological number of the second and third bands $\mathrm{Ch}^{2+3}=-1$. Consequently, these topological excitations are chiral.
Two-dimensional topological spin-orbital supersolid.
In the intermediate regime with moderate $m_{z}$, the $p$-orbital interaction and the Raman-induced SOC are competitive, and lead to an equal condensation at momenta $\widetilde{\bm{k}}_{c}^{n}$ ($n=1\sim 4$) with the condensation parameters satisfying $|\gamma_{\widetilde{\bm{k}}_{c}^{n}}|^{2}=\frac{1}{4}$ and $\beta_{\widetilde{\bm{k}}_{c}^{n},p_{x\downarrow}}=-(-1)^{n}\beta_{\widetilde{\bm{k}}_{c}^{n},p_{y\downarrow}}$. Here the corresponding OAM order, which is given by $\langle L_{\bm{i},z}\rangle=2n_{0}\sum_{n,n^{\prime}=1}^{4}\mathrm{Im}[e^{\mathrm{-i}(\tilde{\bm{k}}_{c}^{n}-\tilde{\bm{k}}_{c}^{n^{\prime}})\cdot\bm{i}}\gamma_{\tilde{\bm{k}}_{c}^{n}}^{\ast}\beta_{\tilde{\bm{k}}_{c}^{n},p_{x\downarrow}}^{\ast}\gamma_{\tilde{\bm{k}}_{c}^{n^{\prime}}}\beta_{\tilde{\bm{k}}_{c}^{n^{\prime}},p_{y\downarrow}}]$, exhibits a staggered positive and negative clusters distribution shown in Fig.4(a), with the periods of $A(B)$-type clusters being $\pi/k_{c}$ in $x$ and $y$ directions. The staggered clusters pattern satisfies $\sum_{\mathrm{Cluster}}\langle L_{\bm{i},z}\rangle\neq 0$ but $\sum_{\mathrm{Global}}\langle L_{\bm{i},z}\rangle=0$, which means that the time-reversal symmetry is broken within each cluster but recovered by averaging for the global system. The supersolid phase exhibits nontrivial topology at each OAM cluster, as characterized by the chiral and anti-chiral edge states of topological excitations on the boundaries of the OAM clusters A ($\langle L_{\bm{i},z}\rangle>0$) and B ($\langle L_{\bm{i},z}\rangle<0$), respectively [see Fig.4(b)] SM . The chiral and anti-chiral edge states are protected by the opposite Chern numbers of the bulk topological excitations in the A and B clusters, and they are robust under impurities or disorders.
The present supersolid phase embodies rich and exotic features beyond the OAM order. In particular, we investigate the real-space spin texture $\langle\bm{\sigma}_{i}\rangle=(\langle\sigma_{i,x}\rangle,\langle\sigma_{i,y}\rangle,\langle\sigma_{i,z}\rangle)$ of the condensate, with the three components given by $(\langle b_{i,s,\uparrow}^{\dagger}\sum_{\nu}b_{i,p_{\nu},\downarrow}+\mathrm{H.c.}\rangle,\langle-\mathrm{i}b_{i,s,\uparrow}^{\dagger}\sum_{\nu}b_{i,p_{\nu},\downarrow}+\mathrm{H.c.}\rangle,\langle n_{i,s,\uparrow}-\sum_{\nu}n_{i,p_{\nu},\downarrow}\rangle)$ with $\nu=x,y$. Here, the spin texture in $x$--$y$ plane $(\langle\sigma_{i,x}\rangle,\langle\sigma_{i,y}\rangle)$ given in Fig.4(c) shows a topological vortex lattice structure characterized by vortex ($B$) and anti-vortex ($A$) distributions, corresponding to the negative (B) and positive (A) OAM cluster centers, respectively. Further, the spin component along $z$ direction $\langle\sigma_{iz}\rangle$ in Fig.4(d) exhibits another staggered clusters pattern, with a period of $1/\sqrt{2}$ that of OAM patterns. From the three spin components the combined whole spin texture $\langle\bm{\sigma}_{i}\rangle$ forms a topological skyrmion lattice structure whose periodicity matches the OAM pattern. With the nontrivial topology encoded in the real space and in the momentum space (for quasiparticles), the superfluidity of $|G\rangle$ renders an unprecedented 2D topological SOSS phase breaking both lattice translation symmetry and $U(1)$ symmetry, as predicted in the present orbital optical Raman lattice.
Conclusion.
[…]
Acknowledgments.
We thank Xin-Chi Zhou, Ting-Fung Jeffrey Poon, and Ye-Bing Zhang for their valuable discussions. This work was supported by National Key Research and Development Program of China (No. 2021YFA1400900), the National Natural Science Foundation of China (Grants No. 12104205, No. 12261160368, and No. 11921005), and the Innovation Program for Quantum Science and Technology (Grant No. 2021ZD0302000) (Z.H.H., K.H.M., B.Z.W., X.J.L.), and by U.S. AFOSR Grant No. FA9550-23-1-0598 (W.V.L.).